OpenStudy (anonymous):

if x/5=10, then 5(x/5)=5(10) this example of which property?

terenzreignz (terenzreignz):

If x = y, then y may replace x in any equation involving it... this is called the substitution property of equality.

OpenStudy (anonymous):

i don't see that in the options

terenzreignz (terenzreignz):

The options are...?

OpenStudy (anonymous):

Reflexive Property Commutative Property for Addition Addition Property of Equality Transitive Property of Equality Multiplication Property of Equality

terenzreignz (terenzreignz):

oh, right, I can see that now :)

OpenStudy (anonymous):

sorry for not including the options before

terenzreignz (terenzreignz):

Well, when we changed \[\Large \frac{x}5 = 10\]to this:\[\Large \color{red}5\left(\frac{x}5\right)=\color{red}5(10)\] What did we do?

OpenStudy (anonymous):

you added the five for each parenthesis?

terenzreignz (terenzreignz):

added? Or multiplied? ;)

OpenStudy (anonymous):

multiplied

terenzreignz (terenzreignz):

So... what does your heart tell you? lol

OpenStudy (anonymous):

nothing it doesn't speak lol

OpenStudy (anonymous):

only beats

terenzreignz (terenzreignz):

haha.... well, you'd best figure it out from here, it's all but obvious now :P

OpenStudy (anonymous):

Transitive?

terenzreignz (terenzreignz):

Is this a quiz? lol

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i wish tho

terenzreignz (terenzreignz):

Oh, then, you can refer to your textbook and review your properties of equalities :)

OpenStudy (anonymous):

was the correct?

OpenStudy (anonymous):

opening the book gives me an automatic headache

OpenStudy (anonymous):

not a math person

OpenStudy (anonymous):

but is it transitive?

terenzreignz (terenzreignz):

What IS the transitive property?

OpenStudy (anonymous):

hold on let me open the book

OpenStudy (anonymous):

Thanks for the help ill try and do it myself

terenzreignz (terenzreignz):

And get yourself an aspirin :P